OpenStudy (anonymous):

Use a determinant to find the area of the triangle with the given vertices (0, 0), (3, 6), (6, −2)

OpenStudy (ybarrap):

Let $$ A=3\hat i+6\hat j\\ B=6\hat i-2\hat j $$ Find \(|A\times B|\) (i.e. the determinant of the cross product): $$ det\begin{bmatrix} \hat{i}& \hat{j} & \hat{k}\\ 3& 6& 0\\ 6& -2 &0 \end{bmatrix}\\ =|\hat i(0)-\hat j(0)+\hat k(-6-36)|\\ =|-42\hat k|\\ =42 $$

OpenStudy (anonymous):

would I divide it by 2 ?

OpenStudy (ybarrap):

yes, you are correct!